# Adapting Large Multimodal Models to Distribution Shifts: The Role of In-Context Learning

### v1

@@ 4.1 Main Results
Hosp0 Hosp1 Hosp2 Hosp3 Hosp4 Acc Sou Tar Acc
@@
Table 2: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on HAM10000 and NIH_Chest datasets. Mean and standard deviation values are calculated over three independent runs, with the best results highlighted in bold.
@@
4.2 Ablation Study
To assess the impact of enhanced invariance on model adaptability, we conduct an ablation study focusing on the Gemini model’s one-shot performance. This study compares three configurations: a baseline using TopKNearestPR, the baseline only with $\mathcal{L}_{\text{cls}}$, and the full InvariantSelectPR that incorporates both $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. Table 3 displays incremental performance gains across datasets with the successive additions of $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. The addition of $\mathcal{L}_{\text{cls}}$ alone leads to a modest increase in performance by $1.77\%$. However, when incorporated with $\mathcal{L}_{\text{CCI}}$, there is a more substantial performance boost of $4.01\%$, confirming the effectiveness of CCI loss in improving the models’ adaptability.
Table 3: Ablation study on loss terms. The baseline is TopKNearestPR.
Configurations Camelyon17 COVID HAM10000 NIH_Chest Average
@@ 4.2 Ablation Study
4.3 In-depth Analysis
4.3.1 ICL vs. Traditional Supervised Finetuning (SFT)
Recent advances in LMMs demonstrate their impressive zero-shot generalization, often outperforming fine-tuned models in natural distribution shifts (Han et al., 2024). This raises questions about whether LMMs with in-context learning, can exceed fine-tuned model performance in scientific datasets, traditionally reliant on domain-specific fine-tuning. We assess our ICL strategy against traditional SFT to explore this. For SFT, we focus on maintaining domain invariance and targeting the class prediction objective, similar to Eq. (5). We fine-tune a CLIP-ViT on source domain data and then apply it to predict outcomes on target examples. This comparison directly measures the effectiveness of ICL versus conventional SFT. For InvariantSelectPR, we choose ICL examples ranging from one to seven and report the best accuracy. Figure 4 displays the comparative performance across four datasets. SFT demonstrates a substantial improvement over zero-shot capabilities with accuracy improvements of 32.4%, and 12.3% on Camelyon17 and HAM10000, but underperforms 1.5% and 5.3% on NIH_Chest and COVID. In contrast, our proposed InvariantSelectPR with few-shot examples, consistently exceeds SFT, with gains of 1.4%, 4.4%, 1.6%, and 6.4% in the same datasets.
@@ 4.3 In-depth Analysis
4.3.2 Backbone Evaluation
We evaluate the impact of different vision encoder backbones on the effectiveness of our InvariantSelectPR method compared to TopKNearestPR. This includes ViT models pretrained on ImageNet-21K and trained with the self-supervised DINO method on ImageNet-1K. Table 4 shows that InvariantSelectPR consistently outperforms TopKNearestPR across datasets, with an average accuracy of 42.9% versus 39.7%. This underscores the limitations of relying solely on pretrained visual similarity for selecting meaningful in-context examples. InvariantSelectPR also demonstrates more consistent performance, with less deviation from mean accuracy (under 1%) compared to TopKNearestPR (nearly 2%). An important observation is the enhanced performance of backbones utilizing self-supervised or contrastive learning methods, supporting the effectiveness of self-supervised learning in capturing generalizable features that contribute to more robust ICL performance, as suggested in studies (Chen et al., 2020; Radford et al., 2021; Wang et al., 2023a).
Table 4: Performance comparison of ICL methods with different vision encoder backbones.
Methods Backbones Camelyon17 COVID HAM10000 NIH_Chest Average
@@ 4.3.2 Backbone Evaluation
4.3.3 ICL Examples with Various Shots
@@
4.3.4 Evaluation Across Different LMMs
The open-source LMMs like IDEFICS (Laurençon et al., 2024) and OpenFlamingo (Awadalla et al., 2023) primarily focus on text and ignore the input signal of images (Bertini Baldassini et al., 2024). Furthermore, these LMMs lack instruction-following ability to choose the response from the answer list. Thus, we use three proprietary LMMs in this comparative analysis: Gemini Pro, GPT-4V, and Claude 3 Opus (Anthropic, 2024). Due to the high computational demands and associated costs of GPT-4V and Claude 3 Opus, we limit our testing to a single dataset, HAM10000, and perform a one-shot evaluation. Figure 6 demonstrates that InvariantSelectPR consistently outperforms other methods across all three LMMs. This method not only exceeds baseline zero-shot performance but also significantly enhances adaptability. Both GPT-4V and Claude 3 Opus exhibit substantial improvements using all ICL methods over their zero-shot capabilities, suggesting that ICL can effectively boost the adaptability of LMMs. This analysis highlights the capacity of InvariantSelectPR to leverage domain-invariant features to enhance LMMs performance under variable conditions.
@@ 4.3.4 Evaluation Across Different LMMs
4.3.5 Distance Metric Evaluation
We examine the effects of employing various distance metrics, including Cosine, Euclidean, and Manhattan, within TopKNearestPR and InvariantSelectPR methods, as illustrated in Table 5. Our findings indicate that the InvariantSelectPR method consistently achieves higher performance than TopKNearestPR across all metrics tested on both the Camelyon17 and HAM10000 datasets.
Table 5: Performance comparison using different distance metrics.
Method Camelyon17 HAM10000
Cosine Euclidean Manhattan Avg Cosine Euclidean Manhattan Avg
TopKNearestPR 61.96 60.77 59.18 60.64 29.84 29.46 30.80 30.03
InvariantSelectPR 63.90 61.09 62.70 62.56 41.56 37.22 38.93 39.24
4.3.6 Computational Efficiency
Table 6 illustrates the trade-off between computational cost and accuracy improvement. While InvariantSelectPR incurs a slightly higher inference time and GPU usage than the zero-shot baseline but offers an 8.60% accuracy improvement. The increased cost is due to the model loading and similarity calculation. InvariantSelectPR’ lower inference time compared to TopKNearestPR is because it loads the vision encoder once per environment instead of for each target sample. Future work will focus on optimizing these steps to reduce inference time while maintaining accuracy gains.
Table 6: Performance comparison of different one-shot ICL methods on the Camelyon17 dataset, in terms of inference time, GPU usage, and accuracy improvement over the zero-shot baseline.
Method Inference Time (s/query) GPU Usage (GB) Accuracy Improvement
@@ 4.3.6 Computational Efficiency
5 Conclusion and Discussion
Concusion. We investigated the efficacy of in-context learning (ICL) to improve the adaptability of LMMs to distribution shifts through our novel ICL approach, InvariantSelectPR. This method not only outperforms standard zero-shot capabilities but also exceeds other methods like RandomPR and TopKNearestPR in handling domain-specific shifts. Evaluations across four datasets confirmed that InvariantSelectPR enhances LMM adaptability by optimally selecting demonstrative examples. Our study offers insights for future work on distribution shifts in foundation models.
Limitations.  Our study is constrained by several factors. Firstly, we confined our analysis to a small selection of benchmark datasets and relied exclusively on commercial and proprietary models, such as Gemini Pro, GPT-4V, and Claude 3 Opus. The limited availability of comprehensive documentation for these models constrains our understanding of their pre-training data, architecture, and inherent biases. This is critical as some broadly used open-source LMMs can not effectively understand multiple images like Flamingo (Alayrac et al., 2022) and simultaneously follow instructions like LLaVA (Liu et al., 2023), necessitating the use of commercial models. Additionally, the substantial financial and computational resources required to access these proprietary models may restrict further validation and analysis. Secondly, our empirical tests involved just 450 samples, which, despite prior research suggesting stability ranging from 180 to 1800 cases (Han et al., 2024), might not reveal scalability issues or subtle biases in larger datasets. Thirdly, the prevalence of numerous domains in healthcare (Yang et al., 2023a) and scientific research (Ji et al., 2022) presents potential challenges in scaling our method across multiple domains.
Boarder Impacts.  This work introduces a novel ICL method to enhance the adaptability of LMMs to distribution shifts, particularly in scientific and healthcare domains. While this method promises improved model accuracy and reliability, its misuse could amplify biases or yield unreliable outputs from LMMs, leading to negative consequences in these critical fields.
Appendix
More contents are put in the appendix, including:
  • A.
    Related Work.
  • B.
    Experimental Details.
  • C.
    t-SNE Visualizations of Visual Features.
Appendix A Related Work
Distribution Shifts.  The literature on distribution shifts categorizes mitigation approaches into two primary strategies: domain adaptation and domain generalization. Domain adaptation techniques, well-established for scenarios where the target domain is known during training, recalibrate models according to the target data’s statistical properties [Ben-David et al., 2006]. These techniques encompass deep transfer learning, which aligns feature distributions between source and target domains [Sun and Saenko, 2016], unsupervised methods that minimize domain discrepancies [Ganin and Lempitsky, 2015], and the use of benchmarks such as Office-Home [Venkateswara et al., 2017] and DomainNet [Peng et al., 2019]. These benchmarks have propelled advances by introducing more complex and diverse scenarios [Han et al., 2022a, b]. In contrast, domain generalization addresses the more daunting challenge of excelling in completely unseen domains. Strategies here include aligning features across multiple source domains [Li et al., 2018b], separating domain-specific from domain-general features [Piratla et al., 2020, Chattopadhyay et al., 2020], employing meta-learning for optimization across various domains [Li et al., 2018a, Balaji et al., 2018], and using data augmentation to mimic domain variability [Volpi et al., 2018, Carlucci et al., 2019]. Recent studies have observed that LLMs have demonstrated exceptional adaptability when dealing with natural distribution shifts but cannot handle the distribution shifts in specialized areas such as healthcare and scientific research [Radford et al., 2021, Han et al., 2023]. This observation motivates this paper’s exploration of ICL and the development of new ICL strategies under distribution shifts.
In-Context Learning.  In-context learning (ICL), particularly defined in GPT-3 [Brown et al., 2020], originated in LLMs for natural language processing (NLP) tasks. ICL is a proven effective paradigm that leverages context augmented with a few examples to enable LLMs to make predictions [Dong et al., 2022, Lu et al., 2021, Wei et al., 2022b, Min et al., 2022, Dong et al., 2022, Wei et al., 2022a, Wolf et al., 2023, Wies et al., 2024, Xie et al., 2021]. The choice of these in-context examples critically impacts performance, as evidenced by studies demonstrating that selecting nearest neighbors based on sentence encoders can significantly enhance the few-shot capabilities of models like GPT-3 [Liu et al., 2021, Wu et al., 2024, Mao and Yu, 2024]. While ICL is established in NLP, it is emerging in visual and multimodal LLMs. The study Flamingo [Alayrac et al., 2022] marks the earliest exploration of visual ICL, with subsequent studies validating the importance of example selection in image painting models [Bar et al., 2022, Wang et al., 2023b, Zhang et al., 2024], visual understanding [Balazevic et al., 2024, Fang et al., 2024], and diffusion models [Wang et al., 2024]. Unlike prior work, this paper uniquely focuses on deeply understanding the role of ICL under distribution shifts, taking a first step in this direction.
Appendix B Experimental Details
B.1 Datasets
This section provides detailed information about the datasets used in the experiments, including their statistics, preprocessing steps, and domain splits.
Dataset Statistics.  We present the data statistics in Table 7. Besides, we use the datasets available in https://huggingface.co/datasets/jameszhou-gl/gpt-4v-distribution-shift and evaluate random_1 subset with 450 cases.
Table 7: Detailed statistics of the datasets used in the experiments.
Dataset Category Prediction Task Domain Type # Domains # Classes # Samples Example Image
Camelyon17 Medical Tumor detection Hospital 5 2 450 [Uncaptioned image]
HAM10000 Medical Skin disease classification Hospital 4 7 450 [Uncaptioned image]
NIH_Chest Medical Lung disease diagnosis Hospital 2 15 450 [Uncaptioned image]
COVID Medical Pneumonia type classification Hospital 2 3 450 [Uncaptioned image]
Preprocessing Steps.  We do not adopt additional preprocessing steps for the images, adhering instead to the guidelines provided by each large multimodal model. For instance, the Gemini model accepts PIL.Image.open(img_path) as input without requiring specific image size, normalization, or augmentation [^4] https://ai.google.dev/gemini-api/docs/get-started/python.
B.2 Prompts
We utilize the following basic prompt template in all ICL experiments, refering to [Han et al., 2024].
image_descriptions = [f"Image {i+1} is {image_class}" for i,
                      (desc, image_class) in enumerate(source_images)]
images_description = ". ".join(image_descriptions)
prompt = f"""Given the images, answer the following question, using the
         specified format.
         {images_description}.
         Question: What is the class of the next image?
         Choices: {’, ’.join(class_names)}.
         Please respond with the following format for each image:
         ---BEGIN FORMAT TEMPLATE---
         Answer Choice: [Your Answer Choice Here]
         Confidence Score: [Your Numerical Prediction Confidence Score Here
         From 0 To 1]
         Reasoning: [Your Reasoning Behind This Answer Here]
         ---END FORMAT TEMPLATE---
         Do not deviate from the above format.
         Repeat the format template for the answer.
B.3 Implementation Details
Vision encoder architectures.  We employ the vit_large_patch14_224_clip_laion2b configuration from the timm library [^5] https://github.com/huggingface/pytorch-image-models/blob/main/timm/models/vision_transformer.py, which utilizes the Vision Transformer (ViT) architecture [Dosovitskiy et al., 2020]. This model divides 224x224 pixel images into 14x14 patches and processes them using a transformer encoder, leveraging extensive pretraining on the LAION-2B dataset with the CLIP approach to enhance visual and textual understanding.
Data and Code.  Our code can be found at the supplementary material and data is publicly available in https://huggingface.co/datasets/jameszhou-gl/gpt-4v-distribution-shift.
Appendix C t-SNE Visualizations of Visual Features
@@ B.3 Implementation Details
In this section, we presnet t-SNE [Van der Maaten and Hinton, 2008] visualizations of the visual features to illustrate how class-conditioned contrastive invariance (CCI) contributes to domain invariance and discriminative capabilities. The visualizations are based on the original vision encoder and our fine-tuned vision encoder on three datasets, as shown in Figure 7. The visual features are extracted using both the pretrained and fine-tuned ViT models. The t-SNE plots are created to highlight the clustering behavior of the features from the target domain. By comparing the plots, we can visually assess the impact of the fine-tuning with CCI on the separation of different classes and the compactness of feature clusters.

### v2

@@ IV-B2 In-Context Selection Through Enhanced Invariance
$\rightarrow$ Hosp0 $\rightarrow$ Hosp1 $\rightarrow$ Hosp2 $\rightarrow$ Hosp3 $\rightarrow$ Hosp4 Avg $\rightarrow$ Sou $\rightarrow$ Tar Avg
@@
TABLE III: Performance comparison of three retrieval methods against the zero-shot approach, illustrating accuracy improvements or decreases on HAM10000 and NIH_Chest datasets. The best and second-best results are marked in Red and Blue respectively. Mean and standard deviation values are calculated over three independent runs. $\rightarrow$” denotes the test scenario.
@@
V Experiments
V-A Experimental Setup
V-A1 Datasets Overview
We use four benchmark datasets to explore distribution shifts, particularly emphasizing domain-specific fine-tuning [15]. Camelyon17 [55] features 450,000 patches from breast cancer images across five hospitals. HAM10000 [56] offers dermatoscopic images critical for skin cancer detection. The NIH_Chest dataset [57] includes over 112,000 X-ray images annotated for thoracic diseases. The COVID dataset [58] provides diverse pneumonia detection data, including COVID-19 cases, from various hospitals. We analyze a practical subset, random_1, with 450 samples [^4] Available at https://github.com/jameszhou-gl/gpt-4v-distribution-shift. Figure I presents detailed statistics of the datasets.
V-A2 Implementation Details
We compare three retrieval methods—RandomPR, TopKNearestPR and InvariantSelectPR—against the baseline zero-shot capability. We employ vit_large_patch14_224_clip_laion2b configuration from the timm library, exploring variations in backbone configurations further in §V-D2. The Gemini model is employed as the primary LMM due to its superior zero-shot performance across varied datasets [15] and its stable log-linear improvement in performance with an increasing number of ICL examples, as observed in a concurrent study [59]. Our main results (§V-B) focus on one-shot performance, with additional insights on the impact of different numbers of shots in §V-D3. We also include other leading LMMs, such as GPT-4V and Claude [11], in our extended analysis in §V-D4. We utilize the following basic prompt template in all ICL experiments as shown in Figure 5.
Figure 5: Basic prompt template in all ICL experiments.
V-A3 Training Protocol
[…]
V-B Main Results
In Tables II and III, our analysis of four benchmark datasets provides a detailed examination of how different ICL methods perform under distribution shifts. The zero-shot approach highlights the inherent ability of LMMs to adapt to new domains without retraining. However, the effectiveness of this adaptability varies significantly with different ICL strategies. The RandomPR strategy, which employs a stochastic method for selecting in-context examples, yields inconsistent results. For instance, on the Camelyon17 dataset, it leads to a slight decrease of 0.3% in accuracy, but it largely underperforms on the COVID, HAM10000, and NIH_Chest datasets, with accuracy decreases of 8.2%, 4.2%, and 3.7%, respectively. This highlights the unpredictable performance of RandomPR across different conditions. Conversely, TopKNearestPR uses a pre-trained vision encoder to identify feature similarities for example selection, leading to a 5.74% improvement on the Camelyon17, which demonstrates the benefits of a more targeted approach in example selection. Despite this success, the method sees declines of 3.9% and 2.9% on the HAM10000 and NIH_Chest datasets, respectively, indicating a lack of consistent performance across all test scenarios. The most effective strategy, InvariantSelectPR, consistently outperforms other methods, significantly exceeding the zero-shot baseline across all datasets, especially achieving remarkable gains of 8.3% on HAM10000 and 8.6% on Camelyon17. These results underscore the importance of advanced in-context example selection techniques in adapting LMMs to distribution shifts. Despite notable gains, the improvements with InvariantSelectPR on NIH_Chest and COVID are modest. In Figure 6, the incremental improvements by InvariantSelectPR align with those from fine-tuned encoders, which generally surpass the fine-tuning approach by 1% to 6%. This suggests that when fine-tuning itself is minimally effective, ICL strategies yield limited enhancements. Future research thus focuses on more sophisticated methods to enhance invariance, beyond fundamental domain-invariance in this work.
V-C Ablation Study
To assess the impact of enhanced invariance on model adaptability, we conduct an ablation study focusing on the Gemini model’s one-shot performance. This study compares three configurations: a baseline using TopKNearestPR, the baseline only with $\mathcal{L}_{\text{cls}}$, and the full InvariantSelectPR that incorporates both $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. Table IV displays incremental performance gains across datasets with the successive additions of $\mathcal{L}_{\text{cls}}$ and $\mathcal{L}_{\text{CCI}}$. The addition of $\mathcal{L}_{\text{cls}}$ alone leads to a modest increase in performance by $1.77\%$. However, when incorporated with $\mathcal{L}_{\text{CCI}}$, there is a more substantial performance boost of $4.01\%$, confirming the effectiveness of CCI loss in improving the models’ adaptability.
TABLE IV: Ablation study on loss terms. The baseline is TopKNearestPR.
Configurations CA CO HA NI Avg
@@ V-C Ablation Study
V-D In-depth Analysis
V-D1 ICL vs. Traditional Supervised Finetuning (SFT)
Recent advances in LMMs demonstrate their impressive zero-shot generalization, often outperforming fine-tuned models in natural distribution shifts [15]. This raises questions about whether LMMs with in-context learning, can exceed fine-tuned model performance in scientific datasets, traditionally reliant on domain-specific fine-tuning. We assess our ICL strategy against traditional SFT to explore this. For SFT, we focus on maintaining domain invariance and targeting the class prediction objective, similar to Eq. (5). We fine-tune a CLIP-ViT on source domain data and then apply it to predict outcomes on target examples. This comparison directly measures the effectiveness of ICL versus conventional SFT. For InvariantSelectPR, we choose ICL examples ranging from one to seven and report the best accuracy. Figure 6 displays the comparative performance across four datasets. SFT demonstrates a substantial improvement over zero-shot capabilities with accuracy improvements of 32.4%, and 12.3% on Camelyon17 and HAM10000, but underperforms 1.5% and 5.3% on NIH_Chest and COVID. In contrast, our proposed InvariantSelectPR with few-shot examples, consistently exceeds SFT, with gains of 1.4%, 4.4%, 1.6%, and 6.4% in the same datasets.
@@ V-D In-depth Analysis
V-D2 Backbone Evaluation
We evaluate the impact of different vision encoder backbones on the effectiveness of our InvariantSelectPR method compared to TopKNearestPR. This includes ViT models pretrained on ImageNet-21K and trained with the self-supervised DINO method on ImageNet-1K. Table V shows that InvariantSelectPR consistently outperforms TopKNearestPR across datasets, with an average accuracy of 42.9% versus 39.7%. This underscores the limitations of relying solely on pretrained visual similarity for selecting meaningful in-context examples. InvariantSelectPR also demonstrates more consistent performance, with less deviation from mean accuracy (under 1%) compared to TopKNearestPR (nearly 2%). An important observation is the enhanced performance of backbones utilizing self-supervised or contrastive learning methods, supporting the effectiveness of self-supervised learning in capturing generalizable features that contribute to more robust ICL performance, as suggested in studies [53, 7, 61].
TABLE V: Performance comparison of ICL methods with different vision encoder backbones.
Methods Backbones CA CO HA NI Avg
@@ V-D2 Backbone Evaluation
V-D3 ICL Examples with Various Shots
@@
V-D4 Evaluation Across Different LMMs
The open-source LMMs like IDEFICS [62] and OpenFlamingo [63] primarily focus on text and ignore the input signal of images [64]. Furthermore, these LMMs lack instruction-following ability to choose the response from the answer list. Thus, we use three proprietary LMMs in this comparative analysis: Gemini Pro, GPT-4V, and Claude 3 Opus [65]. Due to the high computational demands and associated costs of GPT-4V and Claude 3 Opus, we limit our testing to a single dataset, HAM10000, and perform a one-shot evaluation. Figure 8 demonstrates that InvariantSelectPR consistently outperforms other methods across all three LMMs. This method not only exceeds baseline zero-shot performance but also significantly enhances adaptability. Both GPT-4V and Claude 3 Opus exhibit substantial improvements using all ICL methods over their zero-shot capabilities, suggesting that ICL can effectively boost the adaptability of LMMs. This analysis highlights the capacity of InvariantSelectPR to leverage domain-invariant features to enhance LMMs performance under variable conditions.
@@ V-D4 Evaluation Across Different LMMs
V-D5 Distance Metric Evaluation
We examine the effects of employing various distance metrics, including Cosine, Euclidean, and Manhattan, within TopKNearestPR and InvariantSelectPR methods, as illustrated in Table VI. Our findings indicate that the InvariantSelectPR method consistently achieves higher performance than TopKNearestPR across all metrics tested on both the Camelyon17 and HAM10000 datasets.
TABLE VI: Performance comparison using different distance metrics.
Method Cosine Euclidean Manhattan Avg
Camelyon17
TopKNearestPR 61.96 60.77 59.18 60.64
InvariantSelectPR 63.90 61.09 62.70 62.56
HAM10000
TopKNearestPR 29.84 29.46 30.80 30.03
InvariantSelectPR 41.56 37.22 38.93 39.24
V-D6 Computational Efficiency
Table VII illustrates the trade-off between computational cost and accuracy improvement. While InvariantSelectPR incurs a slightly higher inference time and GPU usage than the zero-shot baseline but offers an 8.60% accuracy improvement. The increased cost is due to the model loading and similarity calculation. InvariantSelectPR’ lower inference time compared to TopKNearestPR is because it loads the vision encoder once per environment instead of for each target sample. Future work will focus on optimizing these steps to reduce inference time while maintaining accuracy gains.
TABLE VII: Performance comparison of different one-shot ICL methods on the Camelyon17 dataset, in terms of inference time, GPU usage, and accuracy improvement over the zero-shot baseline.
Method Time (s/query) GPU (GB) Acc Gains
@@ V-D6 Computational Efficiency
V-D7 t-SNE Visualizations of Visual Features
@@
In this section, we presnet t-SNE [66] visualizations of the visual features to illustrate how class-conditioned contrastive invariance (CCI) contributes to domain invariance and discriminative capabilities. The visualizations are based on the original vision encoder and our fine-tuned vision encoder on three datasets, as shown in Figure 9. The visual features are extracted using both the pretrained and fine-tuned ViT models. The t-SNE plots are created to highlight the clustering behavior of the features from the target domain. By comparing the plots, we can visually assess the impact of the fine-tuning with CCI on the separation of different classes and the compactness of feature clusters.
@@
VI Conclusion
We investigated the efficacy of in-context learning (ICL) to improve the adaptability of LMMs to distribution shifts through our novel ICL approach, InvariantSelectPR. This method not only outperforms standard zero-shot capabilities but also exceeds other methods like RandomPR and TopKNearestPR in handling domain-specific shifts. Evaluations across four datasets confirmed that InvariantSelectPR enhances LMM adaptability by optimally selecting demonstrative examples. Our study offers insights for future work on distribution shifts in foundation models.
Our study has several limitations for further improvements. Firstly, we confined our analysis to a small selection of benchmark datasets and relied exclusively on commercial and proprietary models, such as Gemini Pro, GPT-4V, and Claude 3 Opus. The limited availability of comprehensive documentation for these models constrains our understanding of their pre-training data, architecture, and inherent biases. This is critical as some broadly used open-source LMMs can not effectively understand multiple images like Flamingo [42] and simultaneously follow instructions like LLaVA [67], necessitating the use of commercial models. Additionally, the substantial financial and computational resources required to access these proprietary models may restrict further validation and analysis. Secondly, our empirical tests involved just 450 samples, which, despite prior research suggesting stability ranging from 180 to 1800 cases [15], might not reveal scalability issues or subtle biases in larger datasets. Thirdly, the prevalence of numerous domains in healthcare [68] and scientific research [69] presents potential challenges in scaling our method.